# Data-Juicer 2.0: Cloud-Scale Adaptive Data Processing for and with Foundation Models

### v2

@@ Abstract
The burgeoning field of foundation models necessitates advanced data processing mechanisms capable of harnessing vast and valuable data with various types used by these models. Nevertheless, the current landscape presents unique challenges that traditional data processing frameworks struggle to handle effectively, particularly in handling the complexity of multimodal data. In response, we present Data-Juicer 2.0, a data processing system backed by 100+ data processing operators spanning text, image, video, and audio modalities, supporting more critical tasks including data analysis, synthesis, annotation, and foundation model post-training. With seamless compatibility and dedicated optimization for popular dataset hubs like Hugging Face and computing engines like Ray, it improves upon its predecessor in terms of usability, efficiency, and programmability. It features an easily accessible user interface layer that supports decoupled Python interactions, RESTful APIs, and conversational commands. It contains a new runtime layer optimized for adaptive execution and management across varying dataset scales, processing demands, and computational environments, while hiding unnecessary system details. Extensive empirical evaluations demonstrate Data-Juicer 2.0’s remarkable performance and scalability, highlighting its capability to efficiently process TB-level data with 10k+ CPU cores. The system is publicly available and has been widely adopted in diverse research fields and real-world products such as Alibaba Cloud PAI. We actively maintain it and share insights from practical feedback, with the goal of facilitating research and application of next-generation foundation models.
@@ 1 Introduction
Data Processing Challenges for Foundation Models. Modern foundation models require sophisticated data pipelines to handle multimodal inputs and evolving paradigms for pre-training and post-training. While existing frameworks address specific aspects of text processing together2023redpajama ; DolmaToolkit or traditional big data workloads apache-spark , they lack essential capabilities for contemporary multimodal workflows. Three critical gaps emerge:
Multimodal Processing Limitations: Current tools provide inadequate support for cross-modal alignment and semantic-aware transformations crucial for vision-language models li2024mgm . The transition from text-only systems like the inaugural version, Data-Juicer 1.0.  [^1] In this paper, references to Data-Juicer 1.0 pertain to the SIGMOD publication chen2024dj and the open-source codes versioned with the v0.1.2 Based on it, there are over 100 pull requests (PRs) for updating to Data-Juicer 2.0, including about 40k and 8k lines of Python code addition and deletion, respectively. to multimodal processing introduces architectural challenges in handling heterogeneous data types and inter-modal relationships.
@@
Ecosystem Fragmentation: Disjoint APIs across popular frameworks (Hugging Face Datasets lhoest-etal-2021-datasets , Ray Data ray ) force practitioners into suboptimal workflow choices. The lack of unified abstractions hampers portable optimization and cross-platform execution.
Architecture Overview. Data-Juicer 2.0 addresses these challenges with a layered architecture optimized for foundation models (Fig. 1), enabling multimodal data processing at scale:
@@
Interface Layer: Provides multi-level APIs balancing flexibility and accessibility. Low-level Python APIs enable custom workflows, while RESTful endpoints and a visual editor support rapid prototyping. Novel agent integration allows natural language specification of data pipelines.
@@
Community and Applications. We have open-sourced this new system at https://github.com/modelscope/data-juicer, fostering sustained maintenance and engagement through practical events qin2024synergy ; dj-competition , such as tutorials, surveys, data competitions, and co-optimization with community like Apache Arrow, Ray, and NVIDIA-Nemo-Curator teams.  [^2] Related PRs and docs can be exemplified in links for Arrow, Ray and Nemo respectively. Data-Juicer 2.0 also facilitates many foundation model researches such as those from Alibaba Tongyi ge2024data ; qin2024federated ; bai2024federated ; jiao2024img ; ling2025diversity ; xu2025mindgym ; zhou2024humanvbench , and serves as the operator base for multiple Alibaba Cloud products such as PAI Designer paidesigner2024 , indirectly benefiting hundreds of internal and external customers across various real-world AI businesses. One of them has been running stably for over five months, processing data at a scale exceeding terabytes.
Figure 1: The overview of Data-Juicer 2.0.
@@ 2 Processing Capabilities Beyond Text & Pretraining
Data-Juicer 2.0 extends its predecessor, Data-Juicer 1.0—which focused on pretraining tasks for text-only LLMs—to support the processing of multimodal datasets on more tasks for a wide range of foundation models with about 100 new OPs. Images, videos, and audios are all valid inputs of Data-Juicer 2.0, and it’s able to process datasets for post-training tasks, including supervised fine-tuning (SFT), Reinforcement fine-tuning (RFT), and so on. A detailed list of these OPs is placed in Appendix B. For clarity and ease of reference, the OPs are categorized from the following aspects, and their distributions are visually displayed in Fig. 2:
@@
Figure 2: Distribution of new OPs across various dimensions.
@@
Modality Types: The majority of new OPs are concentrated on video/image/audio/text-only processing, with about 20 OPs dedicated to cross-modal data processing. Among these new OPs, almost 3/4 of them are aimed at multimodal data processing, covering images, videos, and audio. There are both single-modality OPs, such as video_motion_score_filter that scores the dynamics of videos, and cross-modal OPs such as phrase_grounding_recall_filter and video_captioning_from_summarizer_mapper, which measure the alignment between different modalities and generate contents from one modality to another, respectively. Detailed showcases are in Appendix C.
@@
Figure 3: Distribution of fine-grained function types for text-only OPs.
Besides, 25 text-only OPs are introduced for post-training tasks. We tag them with fine-grained function types and summarize their distribution in Fig. 3. Among them, there are information extraction and synthesis OPs for curating dialog samples, as well as calibration and refinement OPs for optimizing questions and responses in SFT and RFT tasks. Beyond previous rule-based OPs to assess quality, difficulty, and diversity of datasets in Data-Juicer 1.0, more foundation-models-based OPs emerge to score the datasets from diverse dimensions and analyze the rationale of given scores. This practice is flexibly and useful for vertical and cross-task scenarios. For example, rule-based methods usually struggle to analyze the difficulties of math and financial problems, while this kind of OPs generally perform better with reduced customization and tuning efforts.
Implementation Types: The new OPs include both novel algorithms from Data-Juicer 2.0 and implementations based on state-of-the-art (SOTA) methodologies schuhmann2021laion ; zhu2023multimodal ; gadre2024datacomp from the community. Some OPs offer multiple versions to accommodate varying computational resources, such as CPU-only (e.g., with OpenCV opencv_library ) or GPU-based configurations (e.g., with RAFT teed2020raft ). In the distribution, model-based OPs dominate, as semantic-aware processing often requires advanced models, such as large models for general-purpose understanding and generation via SDXL, GPT, and Qwen. Data processing with foundation models is becoming more popular and helpful.
@@ 3 Towards a More Accessible System
Data-Juicer 2.0 involves substantial work to improve the system accessibility over Data-Juicer 1.0, including a variety of user interfaces and a smooth experience for both beginners and light users.
On the one hand, in terms of user interfaces, in addition to the way to process data with an all-in-one configuration file provided in the previous version, we introduce 4 more user interaction methods. (1) Low-level APIs. We make the underlying implementation more transparent, and expose many extensible interfaces of Data-Juicer-Dataset, Data-Juicer-Operators, such that users can conveniently integrate Data-Juicer in their code. (2) RESTful APIs. For some web server usages, we provide one-click generation of high-performance web APIs capable of automatically discovering, registering, and adapting OP classes and tools. Users can start the server easily and trigger data processing across computing nodes. (3) Web tools. Based on RESTful APIs, Data-Juicer’s capabilities are integrated into Alibaba Cloud’s visual modeling product, PAI-Designer paidesigner2024 , which provides a drag-and-drop UI for users to organize the data processing pipelines and allows users to make use of more products in the Alibaba Cloud ecosystem. (4) Natural language interaction. It’s challenging to know when and how to use each of the 100+ diverse OPs. In response, we adopted AgentScope gao2024agentscope , a multi-agent platform, for low-code integration, using prompt-optimized reasoning and acting (ReAct) Yao-arxiv-2022-ReAct agents to align OP functionalities with our RESTful APIs. Users only need to tell the agents how they want to process the dataset, and the system will automatically handle the job analysis and execution. More details about these interfaces are in Appendix D.
On the other hand, as the 100+ OPs in Data-Juicer 2.0 cover quite diverse multimodal data and training tasks, the complete capabilities require heavy dependencies and take the risk of runtime errors. This may slow down the system installation, initialization, and troubleshooting. We thus prepare a minimal set of requirements and split OPs’ dependencies into subgroups based on their modalities and usage categories. Users only need to install the lite Data-Juicer in a quick way, and the optional dependencies will be lazy-loaded when using specific OPs. Moreover, we maintain an automated unit and regression testing mechanism, ensuring over 85% test coverage. These actions make our system more user-friendly, helping both beginners and light users get started with it.
@@ 4.1 Data-Juicer-Dataset
Unified Execution Abstraction. Data-Juicer 2.0 introduces a Data-Juicer-Dataset class that abstracts heterogeneous computational engines (Hugging Face Dataset, Ray Data, MaxFrame-DataFrame), while preserving native API compatibility through a Facade-pattern design gamma1995design . It provides unified interfaces for standalone/distributed execution modes, enabling transitions between environments while hiding engine-specific complexities. The class supports chainable processing workflows via templated methods with multiple process() call manners. More implementation details about these modes are provided in Appendix E.1 and Listing E.1.
Reliable Data Loading. Data loading is refactored to a more systematic way instead of only specifying the dataset path. The new DatasetBuilder class supports explicit dataset source specification (e.g., local, remote) and customizable configuration for loading datasets. To improve data loading reliability, a novel DataValidator module is introduced to check and validate data sources, schemas, and whether the dataset meets the goal of data processing. For example, post-tuning data should contain dialogs, and image paths should exist in image-captioning datasets.
Token-Aligned Data Schema. Our intermediate schema represents multimodal data through special tokens (e.g., <__dj__image>, where “dj” stands for Data-Juicer) in text fields, with chunk-based alignment using <|__dj__eoc|> separators. This token-centric design supports both simple cross-modal pairs and complex interleaved datasets (MMC4-style zhu2023multimodal ), while preserving positional relationships and reducing redundant computing of shared media files. Core fields (“text”, “query”/“response”) also align with the popular formats of language foundation models, supplemented by bi-directional conversion tools for training ecosystems like LLaMA-Factory zheng2024llamafactory and ModelScope-SWIFT zhao2024swiftascalablelightweightinfrastructure (implementation details are in Appendix E.2).
@@
Moreover, large-scale data processing faces schema/corruption errors that halt pipelines. Data-Juicer 2.0 supports sample-level fault tolerance through unified batched processing signatures and exception handlers that skip invalid batches (Fig. 7 in Appendix F.2), ensuring worst-case progress guarantees. Schema-compatible empty samples also maintain dataset integrity. To mitigate memory bottlenecks, we introduce streaming JSONL loading via patched Arrow backends, and automatic small file splitting aligning with Ray’s block distribution strategy. This reduces network overhead and improves CPU utilization. These optimizations are detailed in Appendix F.
@@ 4.2 Data-Juicer-Operators
Composable Processing Primitives. We extend Data-Juicer 1.0’s five atomic OPs (Formatter/Filter/Mapper/Deduplicator/Selector) with five compositional types: Grouper/Aggregator/FusedOP/ScriptOP/HumanOP. Using Strategy/Decorator patterns gamma1995design , they enable flexible algorithm encapsulation and runtime behavior extension (OP taxonomy in Appendix E.3.1). FusedOP optimizes batch-wise OP fusion (Appendix E.3.2), while ScriptOP integrates custom Python logic. The Grouper takes a Data-Juicer-Dataset as input and groups data samples into batches, which can then be input into the Aggregator for subsequent aggregation. All OPs follow a unified template method run() with automatic parallelism configuration, decoupling execution logic from runtime engines.
@@
We employ a dedicated Adapter class, which uses a probe_small_batch() method to systematically probe and analyze essential information by applying individual OPs on randomly sampled data in runtime. As a result, we enhance Data-Juicer 1.0 ’s greedy OP fusion with adaptive reordering based on estimated OP speeds. Faster OPs precede slower ones within commutativity constraints, optimizing end-to-end latency (validation in Appendix H.2.1). Moreover, using a uniform parallelism granularity across all OPs in a data pipeline can cause OOM issues for some and resource underutilization for others. In Data-Juicer 2.0, with auto-configuration, model-based OPs use GPU/quantization (e.g., vLLM), while I/O-bound OPs use hierarchical parallelism across batched processing, multiprocessing and multithreading, taking the concurrent opportunities between I/O and computation latencies. The implementation details for OP Adapter are in Appendix G.
OP Insight Mining. The combined effects of sequential OPs are not always additive, as validated in chen2024djsandbox . Existing tools like Data-Juicer 1.0 chen2024dj and Falcon refinedweb focus on coarse-grained metrics (e.g., data volume changes via Sankey diagrams) but lack fine-grained analysis.
To formulate better data recipes, Data-Juicer 2.0 tracks dataset statistics (e.g., perplexity) and semantic tags (e.g., image categories) after each OP execution. Built-in Analyzer leverages Filters and modality-specific tagging OPs to generate histograms of statistical distributions and semantic categories (an example is shown in Appendix, Fig. 8). Data-Juicer 2.0 automates metric comparison between consecutive OPs, producing reports that highlight significant lineage-level variations. For example, a sudden text-length reduction after applying a BLIP-2 li2023blip2 image-text matching Filter could indicate noisy captions requiring adjustment. These insights help users systematically evaluate OP impacts, optimize data recipes, and identify unintended correlations.

### v3

@@ Abstract
Foundation models demand advanced data processing for their vast, multimodal datasets. However, traditional frameworks struggle with the unique complexities of multimodal data. In response, we present Data-Juicer 2.0, a data processing system backed by 100+ data processing operators spanning text, image, video, and audio modalities, supporting more critical tasks including data analysis, synthesis, annotation, and foundation model post-training. With seamless compatibility and dedicated optimization for popular dataset hubs like Hugging Face and computing engines like Ray, it improves upon its predecessor in terms of usability, efficiency, and programmability. It features an easily accessible user interface layer that supports decoupled Python interactions, RESTful APIs, and conversational commands. Its new runtime layer offers adaptive execution across diverse scales and environments, abstracting away system complexities. Extensive empirical evaluations demonstrate Data-Juicer 2.0’s remarkable performance and scalability, highlighting its capability to efficiently process TB-level data with 10k+ CPU cores. The system is publicly available and has been widely adopted in diverse research fields and real-world products such as Alibaba Cloud PAI. We actively maintain the system and share practical insights to foster research and applications of next-generation foundation models.
@@ 1 Introduction
Data Processing Challenges for Foundation Models. Foundation models require sophisticated pipelines for multimodal data across evolving paradigms in pre-training and post-training. While existing frameworks address specific aspects of text processing together2023redpajama ; DolmaToolkit or traditional big data workloads apache-spark , they lack essential capabilities for contemporary multimodal workflows. Three critical gaps emerge:
Multimodal Processing Limitations: Current tools provide inadequate support for cross-modal alignment and semantic-aware transformations crucial for vision-language models li2024mgm . The transition from text-only systems like the inaugural version, Data-Juicer 1.0.  [^1] In this paper, references to Data-Juicer 1.0 pertain to the SIGMOD publication chen2024dj and the open-source codes versioned with the v0.1.2. to multimodal processing introduces architectural challenges in handling heterogeneous data types and inter-modal relationships.
@@
Ecosystem Fragmentation: Disjoint APIs across popular frameworks force practitioners into suboptimal workflow choices. The lack of unified abstractions hampers portable optimization and cross-platform execution.
Figure 1: The overview of Data-Juicer 2.0.
Architecture Overview. Data-Juicer 2.0 addresses these challenges with a layered architecture optimized for foundation models (Fig. 1):
@@
Interface Layer: Provides multi-level APIs balancing flexibility and accessibility. Low-level Python APIs enable custom pipelines, while RESTful endpoints and a visual editor support rapid prototyping. Novel agent integration allows natural language specification of pipelines.
@@
Community and Applications. We have open-sourced this new system at https://github.com/modelscope/data-juicer, fostering sustained maintenance and engagement through practical events qin2024synergy ; dj-competition , such as tutorials, surveys, data competitions, and co-optimization with community like Apache Arrow, Ray, and NVIDIA-Nemo-Curator teams. Data-Juicer 2.0 also facilitates many foundation model researches such as those from Alibaba Tongyi ge2024data ; qin2024federated ; bai2024federated ; jiao2024img ; ling2025diversity ; xu2025mindgym ; zhou2024humanvbench ; jiao2025detailmastertexttoimagemodelhandle , and serves as the operator base for multiple Alibaba Cloud products such as PAI Designer paidesigner2024 and DLC-Ray paidlcray , indirectly benefiting hundreds of internal and external customers across various real-world AI businesses. One of them has been running stably for over five months, processing data at a scale exceeding terabytes.
2 Preliminaries and Design Rationale
2.1 Related Work & Core Challenges
While existing systems address big data apache-spark or text-centric model data together2023redpajama ; DolmaToolkit ; Jennings_NeMo-Curator_a_toolkit , a dedicated, open-source framework for multimodal foundation models is lacking. This gap stems from three core challenges rooted in the unique demands of modern AI workflows qin2024synergy ; bai2024survey : 1) Functionality: Processing requires deep semantic understanding and cross-modal alignment (e.g., for video, image, text, audio) liu2023llava ; li2024mgm , moving beyond the structured data focus of traditional systems. 2) Efficiency: Workloads are dominated by simple, per-sample operations (mappers, filters) at massive scale, often involving costly model inference, which contrasts with the complex aggregation queries optimized by conventional engines apache-spark ; dask . 3) Usability: The practitioner ecosystem is centered on Python, demanding native, intuitive interfaces that abstract away backend complexities Paszke-NeurIPS-2019-Pytorch ; lhoest-etal-2021-datasets . These challenges require new system designs, motivating the specific goals of Data-Juicer 2.0.
2.2 System Design Goals in Data-Juicer 2.0
Comprehensive Multimodal Processing. To address functionality gaps, extensive operators are required for collecting, cleansing, annotating, and synthesizing data across modalities like video, image, text, and audio, integrating both perceptual and cognitive information videoworldsimulators2024 ; podell2023sdxl .
Efficient and Optimized Data Flow. To tackle efficiency issues, we aim to accelerate high-frequency basic operators and efficiently manage high-cost semantic operations, while minimizing I/O and data transfer overhead.
Intuitive and Engine-Agnostic Interface. To enhance usability, we aim to protect users from the complexities of underlying execution engines, with easy-to-use Python APIs, graphical interface, and natural language interaction.
Adaptive and Scalable Execution. The system must adapt to diverse computational environments and workloads. It is designed to intelligently orchestrate and optimize data processing across various backends—from local execution to distributed computation on large-scale clusters.
2.3 Key Differences from Prior Systems
While its predecessor, Data-Juicer 1.0, laid a foundation for text data processing, it faced limitations in multimodal support, programmability, and fragmented workflows across different backends. Data-Juicer 2.0 is architected to overcome these issues. Unlike modality-specific toolkits Jennings_NeMo-Curator_a_toolkit ; penedo2024datatrove or general big data systems apache-spark , Data-Juicer 2.0 introduces a composable operator system that generalizes across data types and training tasks while retaining scalability. It distinctively integrates deep learning models as first-class citizens in the data pipeline, emphasizing both computational efficiency and semantic richness. Table 1 summarizes the key advancements from Data-Juicer 1.0 to Data-Juicer 2.0.
Table 1: Comparison of Data-Juicer 1.0 and 2.0.
Feature Data-Juicer 1.0 New in Data-Juicer 2.0
Data Modality Text-only (~50 OPs) + ~100 Image, video, audio OPs
Major OP Types Formatter, Filter, Mapper + Grouper, Aggregator, FusedOP, HumanOP
Deduplicator Standalone-only + Ray-based distributed deduplication
Interaction CLI, Low-level APIs + RESTful APIs, Web UI, NL Interface
Execution Engines HF Datasets, Beam + Tighter integration with Alibaba PAI, Ray
Optimization Greedy OP fusion + Advanced OP reordering, auto resource allocation, batched/parallel execution, etc.
Compute Scale 1,000+ cores 10,000+ cores
Data Scale ~70M samples (TB-level) ~70B samples (PB-level)
Code Contribution 26 PRs, ~34k LoC + 100+ new PRs, ~40k added, ~8k deleted
@@ 3 Processing Capabilities Beyond Text & Pretraining
Data-Juicer 2.0 extends its predecessor to support the processing of multimodal datasets on more tasks for a wide range of foundation models with about 100 new OPs. Images, videos, and audios are all valid inputs of Data-Juicer 2.0, and it’s able to process datasets for post-training tasks, including supervised fine-tuning (SFT), Reinforcement fine-tuning (RFT), and so on. Detailed list of these OPs is in Appendix A. For clarity and ease of reference, the OPs are categorized from the following aspects, and their distributions are visually displayed in Fig. 2:
@@
Figure 2: Distribution of new OPs across various dimensions. The high-resolution versions of these subfigures can be found in Appendix, Figure 15, 16, 17, and 18.
@@
Modality Types: The majority of new OPs are concentrated on video/image/audio/text-only processing, with about 20 OPs dedicated to cross-modal data processing. Among these new OPs, almost 3/4 of them are aimed at multimodal data processing, covering images, videos, and audio. There are both single-modality OPs, such as video_motion_score_filter that scores the dynamics of videos, and cross-modal OPs such as phrase_grounding_recall_filter and video_captioning_from_summarizer_mapper, which measure the alignment between different modalities and generate contents from one modality to another, respectively. Detailed showcases are in Appendix B.
@@
Figure 3: Distribution of fine-grained function types for text-only OPs. A high-resolution version can be found in Appendix, Figure 19.
Besides, 25 text-only OPs are introduced for post-training tasks. We tag them with fine-grained function types and summarize their distribution in Fig. 3. Among them, there are information extraction and synthesis OPs for curating dialog samples, as well as calibration and refinement OPs for optimizing questions and responses in SFT and RFT tasks. Beyond previous rule-based OPs to assess quality, difficulty, and diversity of datasets in Data-Juicer 1.0, more foundation-models-based OPs emerge to score the datasets from diverse dimensions and analyze rationale of given scores. This practice is flexibly and useful for vertical and cross-task scenarios. For example, rule-based methods usually struggle to analyze difficulties of math and financial problems, while this kind of OPs generally perform better with reduced customization and tuning efforts.
Implementation Types: The new OPs include both novel algorithms from Data-Juicer 2.0 and implementations based on SOTA methodologies schuhmann2021laion ; zhu2023multimodal ; gadre2024datacomp from the community. Some OPs offer multiple versions to accommodate varying computational resources, such as CPU-only (e.g., with OpenCV opencv_library ) or GPU-based configurations (e.g., with RAFT teed2020raft ). In the distribution, model-based OPs dominate, as semantic-aware processing often requires advanced models, such as large models for general-purpose understanding and generation via SDXL, GPT, and Qwen. Data processing with foundation models is becoming more popular and helpful.
@@ 4 Towards a More Accessible System
Data-Juicer 2.0 significantly improves accessibility over Data-Juicer 1.0 with multiple new interfaces catering to both novice and expert users.
In terms of user interfaces, in addition to the way to process data with an all-in-one configuration file provided in the previous version, we introduce 4 more user interaction methods. (1) Low-level APIs. We make the underlying implementation more transparent, and expose many extensible interfaces of Data-Juicer-Dataset, Data-Juicer-Operators, such that users can conveniently integrate Data-Juicer in their code. (2) RESTful APIs. For some web server usages, we provide one-click generation of high-performance web APIs capable of automatically discovering, registering, and adapting OP classes and tools. Users can start the server easily and trigger data processing across computing nodes. (3) Web tools. Based on RESTful APIs, Data-Juicer’s capabilities are integrated into Alibaba Cloud’s visual modeling product, PAI-Designer paidesigner2024 , which provides a drag-and-drop UI for users to organize the data processing pipelines and allows users to make use of more products in the Alibaba Cloud ecosystem. (4) Natural language interaction. It’s challenging to know when and how to use each of the 100+ diverse OPs. In response, we adopted AgentScope gao2024agentscope , a multi-agent platform, for low-code integration, using prompt-optimized reasoning and acting (ReAct) Yao-arxiv-2022-ReAct agents to align OP functionalities with our RESTful APIs. Users only need to tell the agents how they want to process the dataset, and the system will automatically handle the job analysis and execution. More details about these interfaces are in Appendix C.
As the 100+ OPs in Data-Juicer 2.0 cover quite diverse multimodal data and training tasks, the complete capabilities require heavy dependencies and take the risk of runtime errors. This may slow down the system installation, initialization, and troubleshooting. We thus prepare a minimal set of requirements and split OPs’ dependencies into subgroups based on their modalities and usage categories. Users only need to install the lite Data-Juicer in a quick way, and the optional dependencies will be lazy-loaded when using specific OPs. Moreover, we maintain an automated unit and regression testing mechanism, ensuring over 85% test coverage. These actions make our system more user-friendly, helping both beginners and light users get started with it.
@@ 5.1 Data-Juicer-Dataset
Unified Execution Abstraction. Data-Juicer 2.0 introduces a Data-Juicer-Dataset class that abstracts heterogeneous computational engines (Hugging Face Dataset, Ray Data, MaxFrame-DataFrame), while preserving native API compatibility through a Facade-pattern design gamma1995design . It provides unified interfaces for standalone/distributed execution modes, enabling transitions between environments while hiding engine-specific complexities. The class supports chainable processing workflows via templated methods with multiple process() call manners. More implementation details about these modes are provided in Appendix D.1 and Listing 8.
Reliable Data Loading. Data loading is now more systematic, moving beyond simple path specifications. The new DatasetBuilder class supports explicit dataset source specification (e.g., local, remote) and customizable configuration for loading datasets. To improve data loading reliability, a novel DataValidator module is introduced to check and validate data sources, schemas, and whether the dataset meets the processing goal. For example, post-tuning data should contain dialogs, and image paths should exist in image-captioning datasets.
Token-Aligned Data Schema. Our intermediate schema represents multimodal data through special tokens (e.g., <__dj__image>, where “dj” stands for Data-Juicer) in text fields, with chunk-based alignment using <|__dj__eoc|> separators. This token-centric design supports both simple cross-modal pairs and complex interleaved datasets (MMC4-style zhu2023multimodal ), while preserving positional relationships and reducing redundant computing of shared media files. Core fields (“text”, “query”/“response”) also align with the popular formats of language foundation models, supplemented by bi-directional conversion tools for training ecosystems like LLaMA-Factory zheng2024llamafactory and ModelScope-SWIFT zhao2024swiftascalablelightweightinfrastructure (implementation details are in Appendix D.2).
@@
Moreover, large-scale data processing faces errors from data corruption or unreliable operators (e.g., malformed LLM outputs), which can halt entire pipelines. Data-Juicer 2.0 enhances resilience at multiple levels. For operator-specific failures, it performs pre-flight validation on LLM responses and employs automatic retries with backoff. More systematically, to overcome the coarse-grained fault tolerance of underlying engines like Ray (which require full restarts), we introduce operator-level checkpointing and fine-grained recovery. This allows pipelines to resume from the last successful stage rather than from scratch, ensuring robust progress even with intermittent task failures. Schema-compatible empty samples are also used to maintain dataset integrity (e.g., see Fig. 11 in Appendix E.2). These optimizations are detailed in Appendix E.
@@ 5.2 Data-Juicer-Operators
Composable Processing Primitives. We extend Data-Juicer 1.0’s five atomic OPs (Formatter/Filter/Mapper/Deduplicator/Selector) with five compositional types: Grouper/Aggregator/FusedOP/ScriptOP/HumanOP. Using Strategy/Decorator patterns gamma1995design , they enable flexible algorithm encapsulation and runtime behavior extension (OP taxonomy in Appendix D.3.1). FusedOP optimizes batch-wise OP fusion (Appendix D.3.2), while ScriptOP integrates custom Python logic. The Grouper takes a Data-Juicer-Dataset as input and groups data samples into batches, which can then be input into the Aggregator for subsequent aggregation. All OPs follow a unified template method run() with automatic parallelism configuration, decoupling execution logic from runtime engines.
@@
We employ a dedicated Adapter class, which uses a probe_small_batch() method to systematically probe and analyze essential information by applying individual OPs on randomly sampled data in runtime. As a result, we enhance Data-Juicer 1.0 ’s greedy OP fusion with adaptive reordering based on estimated OP speeds. Faster OPs precede slower ones within commutativity constraints, optimizing end-to-end latency (validation in Appendix H.2.1). Moreover, using a uniform parallelism granularity across all OPs in a data pipeline can cause OOM issues for some and resource underutilization for others. In Data-Juicer 2.0, with auto-configuration, model-based OPs use GPU/quantization (e.g., vLLM), while I/O-bound OPs use hierarchical parallelism across batched processing, multiprocessing and multithreading, taking the concurrent opportunities between I/O and computation latencies. The implementation details for OP Adapter are in Appendix F.
OP Insight Mining. The combined effects of sequential OPs are not always additive, as validated in chen2024djsandbox . Existing tools like Data-Juicer 1.0 chen2024dj and Falcon refinedweb focus on coarse-grained metrics (e.g., data volume changes via Sankey diagrams) but lack fine-grained analysis. To formulate better data recipes, Data-Juicer 2.0 tracks dataset statistics (e.g., perplexity) and semantic tags (e.g., image categories) after each OP execution. Built-in Analyzer leverages Filters and modality-specific tagging OPs to generate histograms of statistical distributions and semantic categories (an example is shown in Appendix, Fig. 12). Data-Juicer 2.0 automates metric comparison between consecutive OPs, producing reports that highlight significant lineage-level variations. For example, a sudden text-length reduction after applying a BLIP-2 li2023blip2 image-text matching Filter could indicate noisy captions requiring adjustment. These insights help users systematically evaluate OP impacts, optimize data recipes, and identify unintended correlations.